OpenStudy (anonymous):

solveby substitution method 3m+n=13 m-5n=47

OpenStudy (anonymous):

From the second equation, m = 5n + 47 Substitute that into the first one to get 3(5n + 47) + n = 13 Distribute to get 15n + 141 + n = 13 Combine like terms to get 16n +141 = 13 Subtract 141 from both sides to get 16n = -128 Divide by 16 to get n = -8 Then plug -8 into either equation for n and solve for m.

OpenStudy (anonymous):

That is not correct it is suppose to have two not just one

OpenStudy (anonymous):

I was leaving that part to you, but I'll show you that one too. It would be easier to plug n into the second equation like so: m - 5(-8) = 47 Simplify to get m + 40 = 47 Subtract 40 from both sides to get m = 7